# A Comprehensive Survey on Multi-modal Conversational Emotion Recognition with Deep Learning

### v1

@@ 3.1.6. Multi-modal Conversational Emotion Recognition Versus Graph Neural Network
The GNN-based emotion recognition methods (Lin et al., 2022), (Zhang et al., 2023a), (Li et al., 2023), (Ishiwatari et al., 2020), (Shen et al., 2021) inherit the idea of the RNN method, i.e., the contexts should interact and depend on each other (Ghosal et al., 2019). On the basis of RNN, GNNs believe that there is also a relationship of mutual influence between speakers. Therefore, GNNs model the dialogue relationship between speakers through the inherent properties of the graph structure.
@@ 3.2. Definitions and Preliminaries
Definition 3 (Problem definition) For a given multi-modal utterance sequence $U$, the MCER task requires using the utterance context information to determine a deep neural network $F\left({{u_{i}}}\right)$ so that the output emotion label $\hat{y}_{i}$ is as close as possible to the real emotion label ${y_{i}}$, $i\in\left\{{1,...,L}\right\}$. Deep neural networks can solve the optimal parameters by minimizing loss, and its loss is defined as follows:
@@ 3.3.1. Text Feature Extraction
With the rapid development of deep learning technology, word embedding technology has also been widely used to extract text features. Word embedding technology uses a shallow neural network to learn the semantic information of words, and uses Euclidean distance to measure the similarity between words. Unlike traditional one-hot encoding methods, word embedding technology can map high-dimensional sparse feature vectors to low-dimensional dense feature vectors, thus saving a lot of computing resources and solving the problem that one-hot encoding cannot distinguish the semantic gap between words. A commonly used word embedding method is Word2Vec (Chen et al., 2019), which contains two different forms: CBOW (Ghosh and Anwar, 2021) and Skip-gram (Du et al., 2020). CBOW predicts the central word based on the surrounding words, and Skip-gram predicts the surrounding words based on the central word. Although the above methods can capture the semantic similarity between words, they require large datasets for training.
Some recent studies use simple TextCNN (Kim, 2014) and GLOVE (Gan et al., 2022) to extract text features. In addition, large-scale predictive pre-training models such as BERT (Ma et al., 2021) and RoBERTa (Kim et al., 2021) are often used, which capture contextual information through attention mechanisms.
@@ 3.3.2. Video Feature Extraction
Nowadays, most of the visual features are extracted by some advanced neural networks (e.g., CNN (Kattenborn et al., 2021; Wang et al., 2020a), and Transformer (Han et al., 2021; Han et al., 2022), etc) or some open source toolkits (e.g., OKAO Vision (Zhu et al., 2023), OpenFace (Baltrusaitis et al., 2018), CERT (Boopathy et al., 2019) and Facet (Barreiro and Treglown, 2020), etc). OKAO Vision extracts the speaker’s facial expression in each frame of the video, and finally gets the speaker’s smile intensity (0-100) and eye gaze direction. CERT can adaptively extract visual features such as face and head pose. OpenFace 2.0 is capable of detecting facial landmarks, recognizing facial movements, and estimating eye gaze direction. Facet extracts visual features, including facial movements, head poses, HOG features and etc.
@@ 3.3.3. Audio Feature Extraction
Deep learning has also begun to get more and more attention in research in audio feature extraction research. For example, LSTM (Xie et al., 2019) has been widely used to automatically extract acoustic characteristics. Poria et al. (Poria et al., 2017) uses CNN to extract features from the audio, and then enter the extracted audio features into the classifier for emotional classification.
Recently, more and more emotion analysis models (Tzinis et al., 2018; Majumder et al., 2019; Zhong et al., 2019) have started to use open source toolkits such as CONVAREP (Dumpala et al., 2023), openSMILE (Kumar et al., 2022), LibROSA (Suman et al., 2022), and OpenEAR (Schepker et al., 2020) to extract audio features. Specifically, OpenEAR adaptively extracts a set of acoustic features (e.g., prosody, spectrum and sepstral, etc.) and uses Z-normalization to normalize the audio features. The features extracted by openSMILE consist of MFCC, pitch and sound intensity. The LibROSA Speech Toolkit extracts 33 frame-level acoustic features (i.e., 20-dimensional MFCC and CQT) including speaker intonation variations. Similar to other audio feature extraction methods, COVAREP can also be used to extract features such as 12-dimensional MFCC, maximum dispersion scale (MDQ) and Liljencrants-Fant (LF).
@@ 4.1. Context-free Modeling
These are mostly pioneering works on conversational emotion recognition. Context-free modeling methods aim to learn a feature representation for each sentence, which does not exploit the contextual information of the sentence (Zhang et al., 2020), (Seng et al., 2016), (Lotfian and Busso, 2019). For example, some traditional machine methods (e.g., SVMs (Rozgić et al., 2012; Lin and Wei, 2005), and decision trees (Cichosz and Slot, 2007; Lee et al., 2011), etc) is used to extracts the feature representation of each sentence, and utilize the extracted sentence features to complete emotion classification. The above process assumes that each sentence is independent and does not influence each other. We introduce several common context-free modeling methods based on feature fusion below.
@@ 4.1.1. Add
The early fusion method based on addition operation obtains the final emotional feature representation by weigh-ted summation of different modality features (Deng and Ren, 2021). This fusion method is simple to operate and requires only a small amount of calculation. However, its shortcomings are also obvious. It cannot model the context information in a fine-grained manner, and the information that can be utilized is limited. The formula for implementing the context-free modeling method using the additive approach is defined as follows:
@@
where $h_{e}$ represents the fused emotional vectors, $x^{t},x^{a},x^{v}$ represent the text, audio, and video vectors, respectively.
@@ 4.1.2. Concatenation
where $Concat\left(\cdot\right)$ represents concatenation operation.
4.1.3. Tri-modal Hidden Markov Model
Morency et al. used text, video, and audio features for the task of trimodal emotion analysis, and designed a model to extract useful information in different modal features (Morency et al., 2011). After extracting multi-modal features, the three modal features are connected and input into a Hidden Markov Chain (HMM) classifier (Morency et al., 2011) to learn the emotional state of the input signal. HMM believes that the state of the current moment is only related to the information of the previous moment, which makes the model unable to use the context information of the utterance. The formula of HMM is defined as follows:
×$\times P(\lambda_{i}^{a},\lambda_{j}^{v},\lambda_{k}^{t}|x^{a},x^{v},x^{t})$
(6) $P(w|\lambda^{a}_{i},\lambda^{v}_{j},\lambda^{t}_{k},x^{a},x^{v},x^{t})$
×$\times P(\hat{w}_{b}|\lambda_{i}^{a},\lambda_{j}^{v},\lambda_{k}^{t},x^{a},x^{v},x^{t})$
where $P(\hat{w}_{b}|\lambda_{i}^{a},\lambda_{j}^{v},\lambda_{k}^{t},x^{a},x^{v},x^{t})$ represents the probability of predicted label.
@@ 4.1.4. SVM
where $sign(x>0)=1,sign(x=0)=0,sign(x<0)=1$, $\alpha_{i}^{*},b^{*}$ represents the learnable parameters, $\exp\left(-\frac{\|x-z\|^{2}}{2\sigma^{2}}\right)$ represents the kernel function, $N$ is the number of the samples.
@@ 4.1.5. Multiple Kernel Learning
After preprocessing the features of three different modalities, Poria et al. (Poria et al., 2015) constructed two different feature selectors to achieve feature dimensionality reduction. One of the feature selectors is based on circular correlated feature subset selection (CFS), and the other is based on principal component analysis (PCA). The above two feature selectors can not only eliminate redundant information and noise information, but also improve the running speed of the model. After feature selection and dimensionality reduction, the researchers spliced and merged the processed feature vectors and trained a classifier using a multi-kernel learning (MKL) algorithm (Poria et al., 2015). Based on the previous research work, the authors further propose the Convolutional Recurrent Multi-kernel Learning (CRMKL) (Poria et al., 2016) model. CRNKL uses a convolutional recurrent neural network for emotion detection, which can extract contextual information. The formula of MKL is defined as follows:
@@
where $y_{i}$ is the true label, $\alpha,\beta$ are the learnable parameters, $M$ is the feature dimension.
@@ 4.1.6. Select-Additive Learning CNN
CNN is a classic neural network in visual tasks and cannot be directly used for emotion recognition. In order to solve this problem, Kim et al. (Kim, 2014) proposed the TextCNN model, and its overall process is shown in Fig. 5. In order to perform multi-modal emotion recognition, Wang et al. (Wang et al., 2017) proposed the SAL-CNN model, which first uses multi-modal data to fully train the CNN, and then uses Select-Additive Learning (SAL) to improve its versatility and prevent the model from overfitting during training. The SAL method consists of two phases (i.e., selection and addition). In the selection phase, SAL preserves important features and removes noisy information from the latent feature representations learned from neurons. In the addition phase, SAL improves the model’s noise immunity by adding Gaussian noise to the feature representation. The SAL method improves the generalization performance of deep fusion models.
@@ 4.1.7. Tensor Fusion Network
The tensor-based feature fusion method mainly calculates the tensor product of different modal feature representations through Cartesian product to obtain the fused tensor representation (Pandey et al., 2022). Therefore, the above methods need to first map the input multi-modal feature representation into a high-dimensional space, and then map it back to a low-dimensional tensor space for emotion representation. Tensor-based methods are able to capture important high-order interaction information across time, space, and modality. However, the computational complexity of tensor methods is very high and grows exponentially, and there is no fine-grained semantic information interaction between modalities. Zadeh et al. (Zadeh et al., 2017) proposed the multi-modal tensor fusion network TFN. TFN adopts the method of tensor fusion, which can simulate the interaction process between the three modalities of text, audio and video, and effectively fuse multi-modal features. Although TFN can effectively model information interaction within and between modalities, the model complexity of the TFN method is related to the dimensionality of multi-modal features and grows exponentially. The formula of TFN is defined as follows:
@@
(12) $\mathbf{x}^{m}=\begin{bmatrix}\mathbf{x}^{l}\\ 1\end{bmatrix}\otimes\begin{bmatrix}\mathbf{x}^{v}\\ 1\end{bmatrix}\otimes\begin{bmatrix}\mathbf{x}^{a}\\ 1\end{bmatrix}$
where $\otimes$ represents the outer product, $x^{m}$ represents fused vectors.
@@ 4.1.8. Low-rank Tensor Fusion Network
On the basis of TFN, in order to more efficiently fuse multi-modal data, Liu et al. (Liu et al., 2018a) proposed a low-rank tensor method LFM to achieve dimensionality reduction of multi-modal features, so as to improve the fusion efficiency of multi-modal features as shown in Fig. 6. LFM has achieved high performance on many different tasks.
@@
where $\mathbf{w}_{a},\mathbf{w}_{v},\mathbf{w}_{t}$ represents the decomposed low-rank learnable tensor.

### v2

@@ 3.1.6. Multi-modal Conversational Emotion Recognition Versus Graph Neural Network
The GNN-based emotion recognition methods (Lin et al., 2022; Zhang et al., 2023a; Li et al., 2023; Ishiwatari et al., 2020; Shen et al., 2021) inherit the idea of the RNN method, i.e., the contexts should interact and depend on each other (Ghosal et al., 2019). On the basis of RNN, GNNs believe that there is also a relationship of mutual influence between speakers. Therefore, GNNs model the dialogue relationship between speakers through the inherent properties of the graph structure.
@@ 3.2. Definitions and Preliminaries
Definition 3 (Problem definition) For a given multi-modal utterance sequence $U$, the MCER task requires using the utterance context information to determine a deep neural network $F\left({{u_{i}}}\right)$ so that the output emotion label $\hat{y}_{i}$ is as close as possible to the real emotion label $y_{i}$}, $i\in\left\{{1,...,L}\right\}$. Deep neural networks can solve the optimal parameters by minimizing loss, and its loss is defined as:
@@ 3.3.1. Text Feature Extraction
With the rapid development of deep learning, word embedding has also been widely used to extract text features. Before embedding, raw text data typically undergoes several preprocessing steps to improve the quality and consistency of the input. These steps often include tokenization, lowercasing, punctuation and stop-word removal, and in some cases, lemmatization or stemming. Some studies also perform syntactic or dependency parsing to capture the structural relationships between words, which can further enhance the semantic representation in downstream tasks. Word embedding technology then uses a shallow neural network to learn the semantic information of words, and uses Euclidean distance to measure the similarity between words. Unlike traditional one-hot encoding methods, word embedding technology maps high-dimensional sparse feature vectors to low-dimensional dense vectors. This reduces computational resource usage and addresses the issue of one-hot encoding’s inability to capture the semantic gap between words. A commonly used word embedding method is word to vector (Word2Vec) (Chen et al., 2019), which contains two different forms: continuous bag of words (CBOW) (Ghosh and Anwar, 2021) and Skip-gram (Du et al., 2020). CBOW predicts the central word based on the surrounding words, and Skip-gram predicts the surrounding words based on the central word. Although the above methods can capture the semantic similarity between words, they require large datasets for training.
Some recent studies use TextCNN (Kim, 2014) and global vectors for word representation (GLOVE) (Gan et al., 2022) to extract text features. In addition, large-scale predictive pre-training models such as bidirectional encoder representations from Transformers (BERT) (Ma et al., 2021) and robustly optimized BERT approach (RoBERTa) (Kim et al., 2021) are often used to capture contextual information through attention mechanisms.
@@ 3.3.2. Video Feature Extraction
In most modern multimodal emotion recognition systems, visual preprocessing begins with frame sampling, where video is typically downsampled to a fixed frame rate (e.g., 25 or 30 frames per second) to reduce redundancy and maintain temporal resolution (Kattenborn et al., 2021; Wang et al., 2020a). Each frame is then resized (commonly to 224×224 pixels) and normalized using per-channel mean subtraction and standard deviation scaling (e.g., ImageNet normalization settings) to ensure consistency across inputs. For facial region detection and alignment, face detectors (e.g., dlib or MTCNN) are employed to locate the face in each frame, and affine transformations are applied to align facial landmarks to a canonical pose, improving robustness to head movements and variations in scale or rotation. Various open-source toolkits are used to extract deep visual features from the aligned facial regions. For instance, OpenFace 2.0 (Baltrusaitis et al., 2018) detects 68 facial landmarks, estimates head pose, facial action units (AUs), gaze direction, and eye-blink frequency, which are strongly correlated with affective states. Facet further extracts features such as facial muscle activation, histograms of oriented gradients (HOG), emotion intensity scores, and micro-expressions. These features are typically calculated on a frame-by-frame basis and then aggregated over time using statistical functions (e.g., mean, standard deviation, max) or temporal models such as LSTMs. OKAO Vision is a commercial toolkit capable of estimating smile intensity (ranging from 0 to 100) and eye gaze orientation, while CERT adaptively captures head pose dynamics and subtle expressions over short temporal segments. These visual feature extractors are configured with default or task-optimized parameters depending on the specific dataset conditions (e.g., lighting, camera angle, facial occlusion).
@@ 3.3.3. Audio Feature Extraction
Deep learning has increasingly attracted attention in the field of audio feature extraction, enabling automatic modeling of acoustic patterns associated with human emotion. For instance, LSTM networks (Xie et al., 2019) have been widely applied to model temporal dependencies in speech, while Poria et al. (Poria et al., 2017) used convolutional neural networks (CNN) to extract local patterns from audio signals, followed by feeding the extracted features into emotion classification models.
In recent years, an increasing number of emotion recognition models (Tzinis et al., 2018; Majumder et al., 2019; Zhong et al., 2019) have adopted open-source toolkits for systematic and standardized audio feature extraction. Commonly used tools include COVAREP (Dumpala et al., 2023), openSMILE (Kumar et al., 2022), LibROSA (Suman et al., 2022), and OpenEAR (Schepker et al., 2020). These toolkits provide frame-level acoustic descriptors based on well-established speech analysis techniques. Specifically, OpenEAR is capable of extracting a comprehensive set of low-level descriptors, such as prosodic (e.g., pitch, energy), spectral, and cepstral features, and applies Z-score normalization to ensure feature comparability across samples. The openSMILE toolkit is often configured with the INTERSPEECH 2010 or eGeMAPS feature set, extracting Mel-Frequency Cepstral Coefficients (MFCC), pitch, zero-crossing rate, voice intensity, and other prosody-based features. Audio signals are typically resampled to 16 kHz mono channel, then segmented using sliding windows (e.g., 25 ms with 10 ms stride) to generate frame-wise features. LibROSA, a widely used Python-based audio analysis library, is used to extract 33 frame-level acoustic features, including 20-dimensional MFCCs, chroma features, and Constant-Q Transform (CQT), which captures tonal energy variations. Similarly, COVAREP provides features such as 12-dimensional MFCC, Maxima Dispersion Quotient (MDQ), Normalized Amplitude Quotient (NAQ), and Liljencrants–Fant (LF) glottal model parameters, which are valuable for capturing subtle vocal tract changes related to emotional state.
@@ 4.1. Context-free Modeling
These are mostly pioneering works on conversational emotion recognition. Context-free modeling methods aim to learn a feature representation for each sentence, which does not exploit the contextual information of the sentence (Zhang et al., 2020; Seng et al., 2016; Lotfian and Busso, 2019). For example, some traditional machine methods (e.g., SVMs (Rozgić et al., 2012; Lin and Wei, 2005), and decision trees (Cichosz and Slot, 2007; Lee et al., 2011), etc) is used to extracts the feature representation of each sentence, and utilize the extracted sentence features to complete emotion classification. The above process assumes that each sentence is independent and does not influence each other. We introduce several common context-free modeling methods based on feature fusion below.
@@ 4.1.1. Add
The early fusion method based on addition operation obtains the final emotional feature representation by weighted summation of different modality features (Deng and Ren, 2021). This fusion method is simple to operate and requires only a small amount of calculation. However, its shortcomings are also obvious. It cannot model the context information in a fine-grained manner, and the information that can be utilized is limited. The formula for implementing the context-free modeling method using the additive approach is defined as follows:
@@
where $h_{e}$ represents the fused emotional vectors, $x^{t},x^{a},x^{v}$ represent the text, audio, and video vectors, respectively. The Add method is essentially the direct accumulation of different modal information in the same semantic space, which retains the weighted contribution of each modality in the corresponding feature dimension. Through summation, the model can automatically adjust the numerical expression of each modal feature during the learning process, making the information complementary and improving the overall representation ability.
The Add method is relatively simple and easy to understand and implement. By directly merging the features of different modalities, the feature information of different modalities can be fully utilized. For modalities with strong complementarity, the Add method can well capture the correlation between them. However, the features of different modalities may have different scales or importances. The Add method does not consider the difference in importance between the features of different modalities, which may lead to information loss or imbalance. Therefore, the Add method can be considered in scenarios where there is strong complementarity between modalities, limited computing resources, or requirements for model complexity.
@@ 4.1.2. Concatenation
where $Concat\left(\cdot\right)$ represents concatenation operation. The concatenation method does not perform any information compression or mapping on the features of each modality, but directly concatenate the original information by dimension, theoretically retaining the complete expression of each modality.
The concatenate method only requires simple vector concatenation operations, without complex parameter learning or model design. The concatenate method avoids possible information loss in early fusion. However, the feature dimension after concatenation is the sum of each modality, which may lead to sparsity and overfitting, especially on small-scale data sets. In addition, high-dimensional features may increase the computational complexity of subsequent models and simple concatenation cannot explicitly model the nonlinear relationship between modalities. Therefore, when the information quality of each modality is high and complementary to each other, concatenation can effectively retain information.
@@ 4.1.3. SVM
where $sign(x>0)=1,sign(x=0)=0,sign(x<0)=1$, $\alpha_{i}^{*},b^{*}$ represents the learnable parameters, $\exp\left(-\frac{\|x-z\|^{2}}{2\sigma^{2}}\right)$ represents the kernel function, $N$ is the number of the samples. SVM uses a nonlinear kernel function to implicitly model the nonlinear interaction information between modalities.
Through the kernel function, SVM can find the best classification hyperplane in the high-dimensional feature space, effectively perform nonlinear classification, and is suitable for high-dimensional data. When there are fewer training samples, SVM has good generalization ability, but it is time-consuming for large-scale data sets.
@@ 4.1.4. Multiple Kernel Learning
After preprocessing the features of three different modalities, Poria et al. (Poria et al., 2015) constructed two different feature selectors to achieve feature dimensionality reduction. One of the feature selectors is based on circular correlated feature subset selection (CFS), and the other is based on principal component analysis (PCA). The above two feature selectors can not only eliminate redundant information and noise information, but also improve the running speed of the model. After feature selection and dimensionality reduction, the researchers spliced and merged the processed feature vectors and trained a classifier using a multi-kernel learning (MKL) algorithm (Poria et al., 2015). Based on the previous research work, the authors further propose the convolutional recurrent multi-kernel learning (CRMKL) (Poria et al., 2016) model. CRNKL uses a convolutional recurrent neural network for emotion detection, which can extract contextual information. The formula of MKL is defined as:
@@
where $y_{i}$ is the true label, $\alpha,\beta$ are the learnable parameters, $M$ is the feature dimension. MKL achieves flexible fusion of multimodal information at the kernel space level through multi-core combination and weight optimization.
The multiple kernel learning method can combine multiple different kernel functions according to different data characteristics, which helps to process complex data structures. It is suitable for processing a variety of heterogeneous data or multimodal data and can combine information from different modalities. However, calculating the combination of multiple kernel functions may result in high computational costs.
@@ 4.1.5. Select-Additive Learning CNN
CNN is a classic neural network in visual tasks and cannot be directly used for emotion recognition. To solve this problem, Kim et al. (Kim, 2014) proposed the TextCNN model, and its overall process is shown in Fig. 6. To perform multi-modal emotion recognition, Wang et al. (Wang et al., 2017) proposed the SAL-CNN model, which first uses multi-modal data to fully train the CNN, and then uses Select-Additive Learning (SAL) to improve its versatility and prevent the model from overfitting during training. The SAL method consists of two phases (i.e., selection and addition). In the selection phase, SAL preserves important features and removes noisy information from the latent feature representations learned from neurons. In the addition phase, SAL improves the model’s noise immunity by adding Gaussian noise to the feature representation. The SAL method improves the generalization performance of deep fusion models.
@@
The CNN model is relatively simple and has a fast training speed. It can effectively extract local features in text, especially when processing long texts. Although CNN handles long texts well, it has poor adaptability to short texts and structured data and may lose some word order information when processing texts with sequential order.
@@ 4.1.6. Tensor Fusion Network
Figure 7. Illustration of the Tensor Fusion Network (TFN) for tri-modal fusion. The feature vectors from the language ($\mathbf{z}^{l}$), acoustic ($\mathbf{z}^{a}$), and visual ($\mathbf{z}^{v}$) modalities are first augmented with a constant term and then combined via tensor outer product. This operation explicitly captures unimodal, bimodal (e.g., $\mathbf{z}^{l}\otimes\mathbf{z}^{a}$, $\mathbf{z}^{a}\otimes\mathbf{z}^{v}$), and trimodal ($\mathbf{z}^{l}\otimes\mathbf{z}^{v}\otimes\mathbf{z}^{a}$) interactions in a structured tensor space.
As shown in Fig. 7, the tensor-based feature fusion method mainly calculates the tensor product of different modal feature representations through Cartesian product to obtain the fused tensor representation (Pandey et al., 2022). Therefore, the above methods need to first map the input multi-modal feature representation into a high-dimensional space, and then map it back to a low-dimensional tensor space for emotion representation. Tensor-based methods are able to capture important high-order interaction information across time, space, and modality. However, the computational complexity of tensor methods is very high and grows exponentially, and there is no fine-grained semantic information interaction between modalities. Zadeh et al. (Zadeh et al., 2017) proposed the multi-modal tensor fusion network (TFN). TFN adopts the method of tensor fusion, which can simulate the interaction process between the three modalities of text, audio and video, and effectively fuse multi-modal features. Although TFN can effectively model information interaction within and between modalities, the model complexity of the TFN method is related to the dimensionality of multi-modal features and grows exponentially. The formula of TFN is defined as follows:
@@
(10) $\mathbf{x}^{m}$ $=\mathbf{x}^{l}\otimes\mathbf{x}^{v}\otimes\mathbf{x}^{a}$
$=\begin{bmatrix}1&\mathbf{z}_{a}^{\top}\\ \mathbf{z}_{v}&\mathbf{z}_{v}\mathbf{z}_{a}^{\top}\end{bmatrix}\begin{bmatrix}\mathbf{z}_{l}&\mathbf{z}_{l}\mathbf{z}_{a}^{\top}\\ \mathbf{z}_{l}\mathbf{z}_{v}^{\top}&\mathbf{z}_{l}\mathbf{z}_{v}^{\top}\mathbf{z}_{a}^{\top}\end{bmatrix}$
where $\otimes$ represents the outer product, $x^{m}$ represents fused vectors. With the help of tensor outer products, the interaction information of all levels can be systematically preserved.
Tensor fusion methods use tensor decomposition and high-dimensional fusion techniques to map multimodal information into a unified space and capture high-order relationships between modalities through tensor operations. However, the computational overhead of processing tensor operations is high, which may lead to computational bottlenecks during training.
@@ 4.1.7. Low-rank Tensor Fusion Network
On the basis of TFN, in order to more efficiently fuse multi-modal data, Liu et al. (Liu et al., 2018a) proposed a low-rank tensor fusion (LFM) method to achieve dimensionality reduction of multi-modal features, so as to improve the fusion efficiency of multi-modal features as shown in Fig. 8. LFM has achieved high performance on many different tasks.
@@
where $\mathbf{w}_{a},\mathbf{w}_{v},\mathbf{w}_{t}$ represents the decomposed low-rank learnable tensor. LFM essentially retains the multimodal high-order information expression capability of tensor outer product and explicitly models the joint distribution and deep semantic dependencies of different modalities through low-rank approximation.
Low-rank Tensor Fusion Network reduces the complexity of the model by low-rank decomposition of tensors, which can significantly reduce the computational overhead of tensors and reduce memory requirements. However, although low-rank decomposition can reduce model complexity, it may lose some information.